# Few Batches or Little Memory, But Not Both: Simultaneous Space and Adaptivity Constraints in Stochastic Bandits

### 14 Mar 2026 v1

@@ 4.1.1 An Event-Restricted Change of Measure via $\chi^{2}$-Divergence
Many bandit lower bounds compare two environments that differ only on a single arm $j$ and transfer probability statements between the corresponding transcript laws (see, e.g., Kaufmann et al., 2016; Garivier and Kaufmann, 2016; Lattimore and Szepesvári, 2020). Standard Kullback–Leibler- or total-variation-based arguments are transcript-level: they charge the comparison to all pulls of the perturbed arm. For under-sampling events, this is often fundamentally misaligned. If the event of interest is $\{N_{j}(T)\leq n\}$, then only the first $n$ rewards of arm $j$ can affect that event; anything observed from arm $j$ after the $n$-th pull is irrelevant to it.
@@
Setup.
Write $\mathbb{E}_{0},\mathbb{E}_{1}$ for the corresponding expectations.
The algorithm is a non-anticipating policy: for each $t\geq 1$,
$A_{t}=\rho_{t}\bigl(U,A_{1},R_{1},\dots,A_{t-1},R_{t-1}\bigr)\in[K],$
for some measurable map $\rho_{t}$. The observed reward is
$\mathcal{F}_{t}=\sigma(U,A_{1},R_{1},\dots,A_{t},R_{t}).$
Nothing in the next three results depends on batching or memory constraints. For each integer $n\geq 0$, define the prefix sigma-field
@@
Lemma 7 (Budget-event measurability).
For every event $E\in\mathcal{F}_{T}$ and every integer $n\geq 0$,
$E\cap\{N_{j}(T)\leq n\}\in\mathcal{G}_{n}.$
Proof.
and
By construction, the truncated transcript
and the count $\widetilde{N}_{j}^{(n)}(T)$ are $\mathcal{G}_{n}$-measurable.
$\{N_{j}(T)\leq n\}=\{\widetilde{N}_{j}^{(n)}(T)\leq n\}\in\mathcal{G}_{n}.$
Finally, let
$H_{T}:=(U,A_{1},R_{1},\dots,A_{T},R_{T}).$
$E=\{H_{T}\in\Gamma_{E}\}.$
and the right-hand side is $\mathcal{G}_{n}$-measurable. ∎
@@
$\mathbb{E}_{1}[L_{n}^{2}]=\prod_{\ell=1}^{n}\mathbb{E}_{\xi\sim Q}\!\left[\left(\frac{dP}{dQ}(\xi)\right)^{2}\right]=\bigl(1+\chi^{2}(P\|Q)\bigr)^{n}.$
@@
Apply Lemma 9 and rearrange. ∎
@@
and analogous bounds for other global divergences (Bretagnolle and Huber, 1979; Tsybakov, 2008; Lattimore and Szepesvári, 2020). These quantities pay for all pulls of arm $j$ under one of the two environments.
For under-sampling events such as $\{N_{j}(T)\leq n\}$, this can be much larger than the relevant event-level budget. By contrast, Proposition 8 and Lemma 9 pay only for the first $n$ rewards of arm $j$ that are relevant to the event under consideration, independently of how often arm $j$ may be pulled on typical trajectories under either environment. This event-level alignment is precisely what is needed in the application below to turn regret information under a perturbed instance into a probability-level exploration guarantee.
@@ 4.3 From Information to Batch Complexity
Proof.
$N_{i}^{(b)}:=\sum_{t=t_{b-1}+1}^{t_{b}}\mathbb{I}\{A_{t}=i\},$
and let
$N^{(b)}:=(N_{i}^{(b)})_{i\in[K]}.$
$N^{(b)}=c_{b}(M_{t_{b-1}},U,\tau)\qquad\text{a.s.}$
$Y_{i}=\mathbb{I}\Bigl\{\sum_{b=1}^{B}c_{b,i}(M_{t_{b-1}},U,\tau)\geq n\Bigr\},\qquad i\in[K].$
Since $M_{0}$ is fixed, this proves that $Y$ is a deterministic function of
$(U,\tau,M_{t_{1}},\dots,M_{t_{B-1}}).$
$(U,M_{t_{1}},\dots,M_{t_{B-1}}),$
@@
$H(\mathbf{M})=\sum_{b=1}^{B-1}H\!\bigl(M_{t_{b}}\mid M_{t_{1}},\dots,M_{t_{b-1}}\bigr).\leq\sum_{b=1}^{B-1}H(M_{t_{b}})\leq\sum_{b=1}^{B-1}\log_{2}|\mathcal{M}|\leq(B-1)W\leq BW.$
@@ 5 Upper Bound under Simultaneous Space and Adaptivity Constraints
In this section we present a batched and space-bounded algorithm for the stochastic $K$-armed bandit problem with horizon $T$ and rewards in $[0,1]$. The algorithm uses $O(\log T)$ bits of memory, at most $\widetilde{O}(K)$ batches, and achieves expected regret $\widetilde{O}(\sqrt{KT})$. This nearly matches the lower bound of Section 4 up to logarithmic factors.
@@ 5.1 Algorithm Description
At a high level, our algorithm may be viewed as a space-bounded analogue of a standard batched elimination algorithm (Perchet et al., 2016; Gao et al., 2019; Jin et al., 2021b). The main difficulty is that standard batched elimination algorithms keep a large active set of candidate arms across stages, which is impossible under a strict logarithmic memory budget.
Our solution is to sequentialize this candidate-set maintenance. Instead of storing many potentially good arms simultaneously, the algorithm keeps only one incumbent arm together with a small summary of its performance, and then scans the remaining arms one by one as challengers. Each challenger is compared against the incumbent through a multistage test: if it is confidently worse, it is eliminated; if it survives all stages and appears better, it replaces the incumbent. Because the challenger is evaluated from scratch in each outer iteration, this design trades memory for adaptivity while keeping the regret near minimax.
@@
Readers familiar with batched bandit literature will recognize that this is the standard “minimax”-type grid (Perchet et al., 2016; Gao et al., 2019).
@@
Throughout outer iteration $i$, the variable $a_{i}^{\star}$ denotes the current incumbent, and $\widehat{\mu}_{i}^{\star}$ denotes the stored summary of the arm currently held in $a_{i}^{\star}$. The identity $a_{i,\mathrm{init}}$ stays fixed while the challengers are scanned, even though $a_{i}^{\star}$ may later be replaced.
Next, the algorithm scans challengers
$j\in[K]\setminus\{a_{i,\mathrm{init}}\}.$
For each challenger $j$, the algorithm preallocates $i$ comparison slots indexed by $k=1,\dots,i$, where slot $k$ has length $t_{k}$. If $j$ has not been eliminated, slot $k$ is used to pull $j$ and compute $\widehat{\mu}_{i,k}(j)$ from $t_{k}$ fresh samples. Once $j$ is eliminated, the remaining slots are still executed but are filled by the current incumbent. Thus, for each pair $(i,j)$, the slots are fixed in advance; adaptivity changes only which arm occupies the later slots, not the batch boundaries.
@@
$\widehat{\mu}_{i,k}(j)<\widehat{\mu}_{i}^{\star}-2c(t_{k}),$
we mark $j$ as eliminated and fill all later slots of this $(i,j)$ comparison with the current incumbent. If $j$ survives through level $k=i$ and moreover satisfies
$\widehat{\mu}_{i,i}(j)\geq\widehat{\mu}_{i}^{\star}+2c(t_{i}),$
then $j$ replaces the incumbent and we update the stored incumbent summary to $\widehat{\mu}_{i,i}(j)$. After the main outer-loop schedule $i=1,\dots,L$ is completed, all remaining rounds are assigned to the final incumbent $a_{L}^{\star}$ in one final batch.
Algorithm 1 Simultaneously Batched and Space Constrained Algorithm for standard Multi-Armed Bandit
@@
10:  for each $j\in[K]\setminus\{a_{i,\mathrm{init}}\}$ do
11:   Set $\mathrm{eliminated}\leftarrow\textbf{false}$
12:   for $k=1$ to $i$ do
13:    if $\mathrm{eliminated}$ then
14:     Pull arm $a_{i}^{\star}$ for $t_{k}$ rounds
15:    else
16:     Pull arm $j$ for $t_{k}$ rounds and compute its empirical mean $\widehat{\mu}_{i,k}(j)$
17:     if $\widehat{\mu}_{i,k}(j)<\widehat{\mu}_{i}^{\star}-2c(t_{k})$ then
18:      Set $\mathrm{eliminated}\leftarrow\textbf{true}$
19:     end if
20:     if $k=i$ and $\widehat{\mu}_{i,i}(j)\geq\widehat{\mu}_{i}^{\star}+2c(t_{i})$ then
21:      Set $a_{i}^{\star}\leftarrow j$
22:      Set $\widehat{\mu}_{i}^{\star}\leftarrow\widehat{\mu}_{i,i}(j)$
23:     end if
24:    end if
25:   end for
@@
28: Pull arm $a_{L}^{\star}$ for the final exploitation batch
@@
An outer iteration is not a single batch. Each incumbent-refresh block and each preallocated comparison slot indexed by $(i,j,k)$ is a batch. Because the schedule is deterministic, the number of rounds used before the final exploitation batch is also deterministic, so the length of the last batch is known in advance.
@@ 5.2 Algorithm Analysis
Algorithm 1 uses $O(\log T)$ bits of memory and at most $O\!\big(K\log^{2}\log T\big)$ batches. Moreover, for $\delta=T^{-4}$, its expected regret satisfies
$\mathbb{E}\!\left[\mathrm{Reg}(T)\right]=O\!\left(\sqrt{KT\log T}\cdot\log^{2}\log T\right)=\widetilde{O}(\sqrt{KT}).$
Intuitively, the algorithm stores only one incumbent, one challenger, and a constant number of counters and empirical means, so the memory usage is logarithmic. The outer loop $i=1,\dots,L$ forms the main outer-loop schedule, the middle loop scans arms sequentially to avoid storing a large active set, and the inner loop allocates the prearranged comparison slots for the current challenger. This sequential scan increases the number of batches by a factor of $K$, while the multistage comparison contributes the additional $\log\log T$ factor.
@@
because outer iteration $i$ contains one incumbent-refresh block of length $t_{i}$ and K1K-1 challenger comparisons, each with $i$ preallocated slots of lengths $t_{1},\dots,t_{i}$. Hence the final batch length is exactly $T-N_{\mathrm{main}}$ and is known in advance.
@@
Algorithm 1 uses only $O(\log T)$ bits of memory.
@@
At any moment, the algorithm maintains the current outer index $i$, the current comparison level $k$, the current scheduled length $t_{i}$, the fixed inherited initial arm $a_{i,\mathrm{init}}$, the current incumbent arm $a_{i}^{\star}$, one incumbent summary $\widehat{\mu}_{i}^{\star}$, one challenger summary $\widehat{\mu}_{i,k}(j)$, and the Boolean variable that records whether the challenger has already been eliminated.
All indices and counters are integers bounded by $T$, and therefore each requires only $O(\log T)$ bits. The summaries $\widehat{\mu}_{i}^{\star}$ and $\widehat{\mu}_{i,k}(j)$ are empirical averages of at most $T$ rewards in $[0,1]$. It suffices to store them to inverse-polynomial precision in $T$, which also requires $O(\log T)$ bits. Hence the total persistent memory usage is $O(\log T)$. ∎
@@
$\frac{T}{40}\leq N_{\mathrm{main}}\leq\frac{T}{5}.$
@@
$\frac{4T}{5}\leq T-N_{\mathrm{main}}\leq\frac{39T}{40}.$
@@
$H\triangleq\frac{T}{10K}.$
Ignoring ceilings, the recurrence
$t_{i}=\sqrt{Ht_{i-1}}$
gives
$t_{i}=H^{1-2^{-i}}\qquad(i\geq 0).$
Since $L=\left\lfloor\log_{2}\log_{2}H\right\rfloor$, we have
$2^{-L}\geq\frac{1}{\log_{2}H}\qquad\text{and}\qquad 2^{-L}<\frac{2}{\log_{2}H}.$
@@
$t_{L}=H^{1-2^{-L}}\leq H^{1-\frac{1}{\log_{2}H}}=\frac{H}{2},$
and also
$t_{L}=H^{1-2^{-L}}>H^{1-\frac{2}{\log_{2}H}}=\frac{H}{4}.$
Moreover, for every $i\leq L$,
$\frac{t_{i}}{t_{i-1}}=H^{2^{-i}}\geq H^{2^{-L}}\geq H^{1/\log_{2}H}=2.$
Therefore the sequence grows by at least a factor of $2$, and in particular
$\sum_{k=1}^{i}t_{k}\leq 2t_{i}\qquad\text{for every }i\leq L.$
It follows that
$\sum_{i=1}^{L}\sum_{k=1}^{i}t_{k}\leq 2\sum_{i=1}^{L}t_{i}\leq 4t_{L}.$
@@
$N_{\mathrm{main}}\leq K\sum_{i=1}^{L}\sum_{k=1}^{i}t_{k}\leq 4Kt_{L}\leq 4K\cdot\frac{H}{2}=\frac{T}{5}.$
On the other hand, the last outer iteration alone contributes at least
@@
so $N_{\mathrm{main}}\geq T/40$. Therefore $N_{\mathrm{main}}\in[T/40,T/5]$, and the final batch length is the predetermined quantity $T-N_{\mathrm{main}}\in[4T/5,39T/40]$. ∎
@@
Algorithm 1 uses at most $O\!\big(K\log^{2}\log T\big)$ batches.
@@
In outer iteration $i$, the algorithm contains one incumbent-refresh block and, for each of the K1K-1 challengers, at most $i$ preallocated comparison slots. Thus outer iteration $i$ uses at most 1+(K1)i1+(K-1)i batches. Including the final exploitation batch, the total number of batches is at most
$\sum_{i=1}^{L}\bigl(1+(K-1)i\bigr)+1=O(KL^{2}).$
Since $L=O(\log\log T)$, we have $O(KL^{2})=O\!\big(K\log^{2}\log T\big)$. ∎
It remains to bound the regret. We first state a uniform concentration event for the challenger estimates and the incumbent-refresh estimates.
@@
Let $\delta=T^{-4}$ and let $c(\cdot)$ be as defined in the algorithm description. Then with probability at least $1-\frac{1}{T^{2}}$, the following event $E$ holds:
$E\triangleq\Big\{\forall i,k,j:\ |\widehat{\mu}_{i,k}(j)-\mu_{j}|\leq c(t_{k})\ \text{ and }\ |\widehat{\mu}^{\mathrm{init}}_{i}-\mu_{a_{i-1}^{\star}}|\leq c(t_{i})\Big\}.$

### 30 Mar 2026 v2

@@ 4.1.1 An Event-Restricted Change of Measure via $\chi^{2}$-Divergence
Many bandit lower bounds compare two environments that differ only on a single arm $j$ and transfer probability statements between the corresponding transcript laws [KCG16, GK16, LS20]. Standard Kullback–Leibler- or total-variation-based arguments are transcript-level: they charge the comparison to all pulls of the perturbed arm. For under-sampling events, this is often fundamentally misaligned. If the event of interest is $\{N_{j}(T)\leq n\}$, then only the first $n$ rewards of arm $j$ can affect that event; anything observed from arm $j$ after the $n$-th pull is irrelevant to it.
@@
The full formal setup appears in Appendix A.1. For the statements below, it suffices to consider two adaptive-sampling environments $\mathbb{P}_{0}$ and $\mathbb{P}_{1}$ induced by the same non-anticipating policy and differing only on one arm $j$, whose reward laws are $P$ under $\mathbb{P}_{0}$ and $Q$ under $\mathbb{P}_{1}$. Write
$\mathcal{F}_{t}:=\sigma(U,A_{1},R_{1},\dots,A_{t},R_{t})$
for the transcript filtration, and for each integer $n\geq 0$ let
@@
Lemma 20, proved in Appendix A.1, shows that under a pull budget $N_{j}(T)\leq n$, the event in question is already visible in $\mathcal{G}_{n}$.
@@
$\mathbb{E}_{1}[L_{n}^{2}]:=\prod_{\ell=1}^{n}\mathbb{E}_{\xi\sim Q}\!\left[\left(\frac{dP}{dQ}(\xi)\right)^{2}\right]=\bigl(1+\chi^{2}(P\|Q)\bigr)^{n}.$
@@
Applying Lemma 8 and rearranging gives the desired inequality. ∎
@@
and analogous bounds for other global divergences [BH79, TSY08, LS20]. These quantities pay for all pulls of arm $j$ under one of the two environments.
For under-sampling events such as $\{N_{j}(T)\leq n\}$, this can be much larger than the relevant event-level budget. By contrast, Lemma 8 and Corollary 9 pay only for the first $n$ rewards of arm $j$ that are relevant to the event under consideration, independently of how often arm $j$ may be pulled on typical trajectories under either environment. This event-level alignment is precisely what is needed in the application below to turn regret information under a perturbed instance into a probability-level exploration guarantee.
@@ 4.3 From Information to Batch Complexity
Proof Sketch.
The key point is policy commitment. For each batch $b$, once $(t_{b-1},M_{t_{b-1}},U)$ is fixed, the entire within-batch action sequence is fixed as well. Hence the pull-count vector contributed by batch $b$ is determined by the boundary information available when that batch starts. Summing these batch-level pull counts shows that the full pull-count vector $(N_{1}(T),\dots,N_{K}(T))$, and therefore the thresholded profile $Y$, is determined by $(U,\tau,\mathbf{M})$, where $\tau=(t_{1},\dots,t_{B-1})$ is the realized internal grid.
In the static-grid model, $\tau$ is fixed in advance. In the adaptive-grid model, each new boundary is chosen from the same boundary information, so $\tau$ is itself determined recursively by $(U,\mathbf{M})$. Therefore $Y$ is a deterministic function of $(U,\mathbf{M})$. The formal measurability verification is deferred to Appendix A.2. ∎
@@
$H(\mathbf{M}):=\sum_{b=1}^{B-1}H\!\bigl(M_{t_{b}}\mid M_{t_{1}},\dots,M_{t_{b-1}}\bigr)\leq\sum_{b=1}^{B-1}H(M_{t_{b}})\leq\sum_{b=1}^{B-1}\log_{2}|\mathcal{M}|\leq(B-1)W\leq BW.$
@@ 5 Upper Bound under Simultaneous Space and Adaptivity Constraints
In this section we present a batched and space-bounded algorithm for the stochastic $K$-armed bandit problem with horizon $T$ and rewards in $[0,1]$. Throughout the paper, we use $W$ to denote the persistent-memory budget measured in bits. To avoid overloading this notation, in the upper-bound construction we use an integer block-size parameter $S\in[K]$. The algorithm uses $O(S\log T)$ bits of memory, at most $\widetilde{O}(K/S)$ batches, and achieves expected regret $\widetilde{O}(\sqrt{KT})$.
Equivalently, for any bit budget $W$ with $\Omega(\log T)\leq W\leq O(K\log T)$, choosing $S\asymp W/\log T$ yields an implementation using at most $W$ bits and $\widetilde{O}(K/W)$ batches. Here the $\Omega(\log T)\leq W$ lower bound is needed to represent the relevant counters and empirical means to inverse-polynomial precision in $T$, while the $W\leq O(K\log T)$ upper bound is without loss of generality since standard finite-armed bandit algorithms only need to store per-arm summaries for the $K$ arms. This nearly matches the lower bound of Section 4 up to logarithmic factors.
@@ 5.1 Algorithm Description
At a high level, our algorithm may be viewed as a space-bounded analogue of a standard batched elimination algorithm [PRC+16, GHR+19, JTX+21b]. The main difficulty is that standard batched elimination algorithms keep a large active set of candidate arms across stages, which is impossible under a strict memory budget.
Our solution is to keep only one incumbent together with one block of at most $S$ challengers at a time. Instead of storing all potentially good arms simultaneously, the algorithm scans the remaining arms block by block. Within each block, challengers are compared against a frozen benchmark incumbent through a multistage test: if a challenger is confidently worse, it is deactivated for the rest of the block; if it survives all stages and appears better, it may trigger an incumbent update at the end of the block. Because only one challenger block is kept in memory at a time, this design uses $O(S\log T)$ bits while reducing the number of batches by a factor of $S$ relative to the fully sequential scan.
@@
Readers familiar with batched bandit literature will recognize that this is the standard “minimax”-type grid [PRC+16, GHR+19].
@@
Throughout outer iteration $i$, the variable $a_{i}^{\star}$ denotes the current incumbent, and $\widehat{\mu}_{i}^{\star}$ denotes the stored summary of the arm currently held in $a_{i}^{\star}$. The identity $a_{i,\mathrm{init}}$ stays fixed while the challenger blocks are scanned, even though $a_{i}^{\star}$ may later be replaced.
Next, the algorithm partitions
$K]\setminus\{a_{i,\mathrm{init}}\}$
into
$J\triangleq\left\lceil\frac{K-1}{S}\right\rceil$
ordered blocks
$C_{i,1},\dots,C_{i,J},\qquad|C_{i,j}|\leq S.$
For each block $C_{i,j}$, the algorithm freezes the current incumbent as the benchmark
$\bar{a}_{i,j}\triangleq a_{i}^{\star},\qquad\bar{\mu}_{i,j}\triangleq\widehat{\mu}_{i}^{\star}.$
It then preallocates $i$ comparison slots indexed by $k=1,\dots,i$, where slot $k$ contains $t_{k}$ pulls for each arm-position in the block. If arm $a\in C_{i,j}$ is still active, then slot $k$ is used to pull $a$ and compute $\widehat{\mu}_{i,j,k}(a)$ from $t_{k}$ fresh samples. Once $a$ is deactivated, its reserved pulls in the remaining slots are still executed but are reassigned to the frozen benchmark arm $\bar{a}_{i,j}$. Thus, for each pair $(i,j)$, the slots are fixed in advance; adaptivity changes only which challengers remain active inside the block, not the batch boundaries.
@@
$\widehat{\mu}_{i,j,k}(a)<\bar{\mu}_{i,j}-2c(t_{k}),$
we mark $a$ as inactive and fill all later slots reserved for this arm in block $(i,j)$ with the frozen benchmark arm. Crucially, for a fixed slot $k$, the active/inactive status of every arm-position is already determined at the start of that slot from the outcomes of the previous slots $1,\dots,k-1$. Thus the loop over $a\in C_{i,j}$ inside slot $k$ does not require any within-slot feedback: all pulls assigned to slot $k$ can be scheduled simultaneously in one batch, and only after that batch is completed do we update the active set for slot $k+1$. After level $k=i$, let $\tilde{a}_{i,j}$ be an active arm in $C_{i,j}$ with the largest value of $\widehat{\mu}_{i,j,i}(a)$, if such an arm exists. If $\tilde{a}_{i,j}$ exists and moreover satisfies
$\widehat{\mu}_{i,j,i}(\tilde{a}_{i,j})\geq\bar{\mu}_{i,j}+2c(t_{i}),$
then $\tilde{a}_{i,j}$ replaces the incumbent and we update the stored incumbent summary to $\widehat{\mu}_{i,j,i}(\tilde{a}_{i,j})$. After the main outer-loop schedule $i=1,\dots,L$ is completed, all remaining rounds are assigned to the final incumbent $a_{L}^{\star}$ in one final batch.
Our algorithm is summarized in Algorithm 1.
Algorithm 1 Simultaneously Batched and Space Constrained Algorithm for standard Multi-Armed Bandit with block size $S$
@@
10:  Set $J\leftarrow\left\lceil\frac{K-1}{S}\right\rceil$ and partition $K]\setminus\{a_{i,\mathrm{init}}\}$ into ordered blocks $C_{i,1},\dots,C_{i,J}$ of size at most $S$
11:  for $j=1$ to $J$ do
12:   Set $\bar{a}_{i,j}\leftarrow a_{i}^{\star}$ and $\bar{\mu}_{i,j}\leftarrow\widehat{\mu}_{i}^{\star}$
13:   for each $a\in C_{i,j}$ do
14:    Set $\mathrm{active}(a)\leftarrow\textbf{true}$
15:   end for
16:   for $k=1$ to $i$ do
17:    for each $a\in C_{i,j}$ do
18:     if $\mathrm{active}(a)$ then
19:      Pull arm $a$ for $t_{k}$ rounds and compute its empirical mean $\widehat{\mu}_{i,j,k}(a)$
20:      if $\widehat{\mu}_{i,j,k}(a)<\bar{\mu}_{i,j}-2c(t_{k})$ then
21:       Set $\mathrm{active}(a)\leftarrow\textbf{false}$
22:      end if
23:     else
24:      Pull arm $\bar{a}_{i,j}$ for $t_{k}$ rounds
25:     end if
@@
28:   if there exists some active arm in $C_{i,j}$ then
29:    Let $\tilde{a}_{i,j}\in\arg\max_{a\in C_{i,j}:\,\mathrm{active}(a)}\widehat{\mu}_{i,j,i}(a)$
30:    if $\widehat{\mu}_{i,j,i}(\tilde{a}_{i,j})\geq\bar{\mu}_{i,j}+2c(t_{i})$ then
31:     Set $a_{i}^{\star}\leftarrow\tilde{a}_{i,j}$
32:     Set $\widehat{\mu}_{i}^{\star}\leftarrow\widehat{\mu}_{i,j,i}(\tilde{a}_{i,j})$
33:    end if
34:   end if
35:  end for
36:end for
37: Pull arm $a_{L}^{\star}$ for the final exploitation batch
@@
An outer iteration is not a single batch. Each incumbent-refresh block and each preallocated block-level comparison slot indexed by $(i,j,k)$ is a batch. Because the schedule is deterministic, the number of rounds used before the final exploitation batch is also deterministic, so the length of the last batch is known in advance.
@@ 5.2 Algorithm Analysis
There exist universal constants $c,C_{1},C_{2},C_{3}>0$ such that the following holds. For every integer $K\geq 2$, every block size $S\in[K]$, and every horizon $T\geq cK$, if Algorithm 1 is run with $\delta=T^{-4}$, then it defines a static-grid stochastic bandit algorithm that uses no more than $C_{1}S\log T$ bits of persistent memory, no more than $C_{2}\frac{K}{S}\log^{2}\log T$ batches, and satisfies
$\sup_{\nu}\mathbb{E}_{\nu}[\mathrm{Reg}_{T}]\leq C_{3}\sqrt{KT\log T}\,\log^{2}\log T,$
where the supremum is over all $K$-armed stochastic bandit instances with rewards in $[0,1]$.
Intuitively, the algorithm stores only one incumbent, one block of at most $S$ challengers, and a constant number of counters and empirical means per challenger in the current block, so the memory usage is $O(S\log T)$ bits. The outer loop $i=1,\dots,L$ forms the main outer-loop schedule, the middle loop scans arm-blocks sequentially to avoid storing a large active set, and the inner loop allocates the prearranged comparison slots for the current block. The sequential block scan contributes the factor $K/S$ in the batch complexity, while the multistage comparison contributes the additional $\log\log T$ factor. Hence, for any bit budget $W$ with $\Omega(\log T)\leq W\leq O(K\log T)$, choosing $S\asymp W/\log T$ yields an implementation using at most $W$ bits and $\widetilde{O}(K/W)$ batches.
@@
because outer iteration $i$ contains one incumbent-refresh block of length $t_{i}$ and then scans the remaining K1K-1 arm-positions through challenger blocks, each position receiving $i$ preallocated slots of lengths $t_{1},\dots,t_{i}$. Hence the final batch length is exactly $T-N_{\mathrm{main}}$ and is known in advance.
@@
Algorithm 1 uses only $O(S\log T)$ bits of memory.
@@
At any moment, the algorithm maintains the current outer index $i$, the current block index $j$, the current comparison level $k$, the current scheduled length $t_{i}$, the fixed inherited initial arm $a_{i,\mathrm{init}}$, the current incumbent arm $a_{i}^{\star}$, one incumbent summary $\widehat{\mu}_{i}^{\star}$, the frozen benchmark pair $(\bar{a}_{i,j},\bar{\mu}_{i,j})$ for the current block, the identities of the at most $S$ arms in the current block, one active/inactive flag for each of these arms, and at most $S$ empirical summaries $\widehat{\mu}_{i,j,k}(a)$ for the arms currently processed in the block.
All indices and counters are integers bounded by $T$, and therefore each requires only $O(\log T)$ bits. Each arm identity needs $O(\log K)\leq O(\log T)$ bits, so the current block needs $O(S\log T)$ bits. Each empirical summary is an average of at most $T$ rewards in $[0,1]$, and it suffices to store it to inverse-polynomial precision in $T$, which also requires $O(\log T)$ bits. Since there are only $O(S)$ such summaries and flags alive simultaneously, the total persistent memory usage is $O(S\log T)$. ∎
@@
$\frac{T}{40}\leq N_{\mathrm{main}}\leq\frac{3T}{5}.$
@@
$\frac{2T}{5}\leq T-N_{\mathrm{main}}\leq\frac{39T}{40}.$
@@
$H\triangleq\frac{T}{10K},\qquad s_{i}\triangleq H^{1-2^{-i}}\qquad(i\geq 0).$
We first compare the integer schedule $(t_{i})$ with the smooth proxy $(s_{i})$. We claim that for every $1\leq i\leq L$,
$s_{i}\leq t_{i}\leq 2s_{i}.$ (8)
For the lower bound, use induction:
$t_{i}=\left\lceil\sqrt{Ht_{i-1}}\right\rceil\geq\sqrt{Ht_{i-1}}\geq\sqrt{Hs_{i-1}}=s_{i}.$
For the upper bound, the case $i=1$ is
$t_{1}=\lceil\sqrt{H}\rceil\leq\sqrt{H}+1\leq 2\sqrt{H}=2s_{1}.$
Now suppose $i\geq 2$ and $t_{i-1}\leq 2s_{i-1}$. Since $L\geq 1$ implies $H\geq 4$, we have $s_{i}\geq s_{1}=\sqrt{H}\geq 2$, and therefore
$t_{i}=\left\lceil\sqrt{Ht_{i-1}}\right\rceil\leq\sqrt{Ht_{i-1}}+1\leq\sqrt{2Hs_{i-1}}+1=\sqrt{2}\,s_{i}+1\leq 2s_{i}.$
This proves (8).
Next, we estimate the last schedule level. Since $L=\lfloor\log_{2}\log_{2}H\rfloor$, we have
$\frac{1}{\log_{2}H}\leq 2^{-L}<\frac{2}{\log_{2}H}.$
@@
$\frac{H}{4}=H^{1-\frac{2}{\log_{2}H}}<H^{1-2^{-L}}=s_{L}\leq H^{1-\frac{1}{\log_{2}H}}=\frac{H}{2}.$
Combining this with (8) yields
$\frac{H}{4}<t_{L}\leq H.$ (9)
We now control the partial sums. If $1\leq i\leq L-2$, then
$2^{-i}\geq 2^{-(L-2)}=4\cdot 2^{-L}\geq\frac{4}{\log_{2}H},$
and therefore
$t_{i}\leq 2H^{1-\frac{4}{\log_{2}H}}=\frac{H}{8}\qquad(1\leq i\leq L-2).$
Therefore, for every $1\leq i\leq L-2$,
$t_{i+1}\geq\sqrt{Ht_{i}}\geq 2t_{i}.$
Also, since $t_{L-1}\leq 2s_{L-1}<H/2$, we have
$t_{L}\geq\sqrt{Ht_{L-1}}\geq\sqrt{2}\,t_{L-1}.$
Consequently,
$\sum_{k=1}^{L-1}t_{k}\leq 2t_{L-1},\qquad\sum_{k=1}^{L}t_{k}\leq 2t_{L-1}+t_{L}\leq(1+\sqrt{2})t_{L},$
and thus
$\sum_{i=1}^{L}\sum_{k=1}^{i}t_{k}\leq 2\sum_{i=1}^{L-1}t_{i}+\sum_{k=1}^{L}t_{k}\leq 4t_{L-1}+(1+\sqrt{2})t_{L}\leq(1+3\sqrt{2})t_{L}<6t_{L}.$
@@
$N_{\mathrm{main}}\leq K\sum_{i=1}^{L}\sum_{k=1}^{i}t_{k}<6Kt_{L}\leq 6KH=\frac{3T}{5}.$
For the lower bound, the last outer iteration alone contributes at least
@@
where we used (9). Hence $N_{\mathrm{main}}\geq T/40$. Therefore $N_{\mathrm{main}}\in[T/40,3T/5]$, and the final batch length is the predetermined quantity
$T-N_{\mathrm{main}}\in\left[\frac{2T}{5},\frac{39T}{40}\right].$
@@
Algorithm 1 uses at most $O\!\big(\frac{K}{S}\log^{2}\log T\big)$ batches.
@@
In outer iteration $i$, the algorithm contains one incumbent-refresh block and
$J=\left\lceil\frac{K-1}{S}\right\rceil$
challenger blocks, each contributing at most $i$ preallocated comparison slots. Thus outer iteration $i$ uses at most $1+Ji$ batches. Including the final exploitation batch, the total number of batches is at most
$\sum_{i=1}^{L}\bigl(1+Ji\bigr)+1=O\!\left(\frac{K}{S}L^{2}\right).$
Since $L=O(\log\log T)$, we have $O\!\left(\frac{K}{S}L^{2}\right)=O\!\big(\frac{K}{S}\log^{2}\log T\big)$. ∎
It remains to bound the regret. We first state a uniform concentration event for the block-level challenger estimates and the incumbent-refresh estimates.
@@
For every $(i,j,k,a)$ such that arm $a$ is sampled as a challenger at level $k$. Let $\delta=T^{-4}$ and let $c(\cdot)$ be as defined in the algorithm description. Then with probability at least $1-\frac{1}{T^{2}}$, the following event $E$ holds:
$E\triangleq\Big\{\forall i,j,k,\forall a\in C_{i,j}:\ |\widehat{\mu}_{i,j,k}(a)-\mu_{a}|\leq c(t_{k})\ \text{ and }\ |\widehat{\mu}^{\mathrm{init}}_{i}-\mu_{a_{i-1}^{\star}}|\leq c(t_{i})\Big\}.$